OpenStudy (anonymous):

using de'moivre formula find the indicated power. (3+4i)^3

OpenStudy (irishboy123):

start with : \( (3+4i)^3= 5^3 \left(\dfrac{3}{5}+i \; \dfrac{4}{5} \right)^3 \)